OpenStudy (4meisu):

A company offers it employees a choice of two salary schemes A and B over a period of 10 years. Scheme A = Starting salary $11000 in the 1st year and then an annual increase of $400 per year. Calculate the total (amount of) salary paid over ten years (scheme a)

OpenStudy (anonymous):

128000

OpenStudy (anonymous):

(400x9) + (400x8)...+ (400x1) +(11000x10)

OpenStudy (anonymous):

\[ 400 ( 1 +2 +3 +4+5+6+7+8+9) = 400 \frac {9(9+1)}2= 400(45)=\\18000+ 110000=128000\]